# The collision and merger products of stars do not look alike: A magnetohydrodynamics comparison

### v1

@@ 2.3 Initial conditions
For cells with densities greater than the threshold given above, $K_{1}+K_{2}\approx 1$. This is not true for very low-density cells, since the ‘vacuum’ cells initially have $K_{1}=K_{2}=0$. Thus, $K$ quantifies the mixing through the course of each interaction, with $K>0$ (or $K_{1}>K_{2}$) indicating a higher fraction of the heavier star and $K<0$ (or $K_{1}<K_{2}$) indicating a higher fraction of the lighter star.
@@
Figure 2: Mid-plane snapshots of a stellar collision between a $10\,\mathrm{M}_{\odot}$ star and a $5\,\mathrm{M}_{\odot}$ star when $b=0.25$, $10\,\mathrm{Myr}$ into their MS lifetimes. The three rows illustrate slices, in the collision plane, of densities $\rho$ (top), passive scalars $K$ (center), and magnetic field magnitudes $B$ (bottom). The four columns illustrate these quantities at the times $t$ before collision (leftmost), just before collision (center-left), just after collision (center-right), and after $\sim 60\,t_{\mathrm{dyn},r_{\mathrm{tot}}}$ following collision (rightmost). The core of the collision product is dominated by the initially less massive star, while the outer layers are composed of material from both stars in varying fractions. Furthermore, the amplification of the magnetic field is clearly visible across the panels.
@@
2.3.1 Collisions
We assume the orbits in the collision scenario to be parabolic. This is a reasonable assumption since young stellar clusters are observed to have velocity dispersions $\sigma\sim 10$$20\,\mathrm{km}\,\mathrm{s}^{-1}$ (e.g., M. Gieles et al., 2010), yielding near-parabolic eccentricities [^2] For a hyperbolic orbit, $e=1+r_{\mathrm{p}}v_{\infty}^{2}/GM$, where $r_{\mathrm{p}}$ is the periapsis distance, $v_{\infty}^{2}\sim\sigma$ is the velocity at infinity and $M$ is the total mass. of $|e-1|\sim 10^{-4}$ for our typical stellar masses and radii. It should be mentioned that collisions are not very frequent in young stellar clusters; nonetheless, we consider these parameters for comparative purposes. We define the collision impact parameter as $b=r_{\mathrm{p}}/r_{\mathrm{tot}}$, where $r_{\mathrm{p}}$ is the closest approach distance between the centers of mass of the two stars. In 6 of the 8 collision simulations, we set $b=0.25$. In the other 2 simulations, we set $b=0.50$ to compare the effect of the impact parameter on the collision product. We initially position the relaxed stars $5\,r_{\mathrm{tot}}$ apart prior to the pericenter passage.
2.3.2 Mergers
Unlike collisions, in reality, mergers occur on much longer timescales. For computational feasibility, we speed up the orbit shrinkage phase by adding artificial deceleration terms (similar to R. Pakmor et al., 2013, 2021). The equations below, which are the same as Equation 2 in R. Pakmor et al. (2021) [^3] Note that there is an error in that paper, which is corrected in our equation., quantify the anti-parallel, azimuthal acceleration terms that result in an inspiral at a constant rate $v_{\mathrm{insp}}$:
@@ 2.3.2 Mergers
$\mathbf{a}_{\star 1}=-\dfrac{m_{\star 2}^{2}\mathbf{v}_{\star 1}}{v_{\star 1}^{2}}f(r)$ (4a)
$\mathbf{a}_{\star 2}=-\dfrac{m_{\star 1}^{2}\mathbf{v}_{\star 2}}{v_{\star 2}^{2}}f(r)$ (4b)
Here, $r$ is the instantaneous distance between the two stars, $\mathbf{v}_{\star}$ and $\mathbf{a}_{\star}$ represent the star’s velocity and acceleration vectors, and $v_{\mathrm{insp}}=10\,\mathrm{km}\,\mathrm{s}^{-1}$ is the inspiral velocity we choose. In comparison, typical Keplerian circular velocities for the masses and distances considered are $\sim 500$$600\,\mathrm{km}\,\mathrm{s}^{-1}$. This deviation from the circular velocity results in a slightly eccentric inspiral with $e\approx 0.1$. For a constant value of $v_{\mathrm{insp}}$, stars with a more skewed mass ratio merge quicker (see Section 3 and Figure 7).
We position the relaxed (non-rotating) stars in a circular binary orbit with an initial semimajor axis of $1.5\,r_{\mathrm{tot}}$ and run the simulations with the sped-up inspiral. As a comparison, using the Roche-lobe radius approximation formula from P. P. Eggleton (1983), Roche-lobe filling ensues at $\lesssim 1.31\,r_{\mathrm{tot}}$ for our range of mass ratios (see Table 1).
3 Results
In this Section, we use the shorthand notation $m_{\star 1}$ + $m_{\star 2}$’ to denote an interaction between stars of masses $m_{\star 1}$ and $m_{\star 2}$ (in units of $\mathrm{M}_{\odot}$). We focus on comparing key properties of the collision and merger products, including chemical mixing, rotation, magnetic fields, and outflows. While these quantities are intrinsically at least 2D, we discuss them using their mass-weighted averages over equipressure surfaces as representative values [^4] MESA handles stellar rotation in 1D in a similar manner, with pressure being the independent coordinate.. In addition to the averages, we indicate the spread in each calculated quantity within the equipressure surfaces by reporting the 15th and 85th percentiles. In the calculation of the averages and spreads, we only consider cells that are bound to the center of mass (CoM) of the new star, which we calculate iteratively as follows. We first find the initial values of the CoM position $\mathbf{r}_{\mathrm{CoM}}$ and velocity $\mathbf{v}_{\mathrm{CoM}}$ of all cells. We then calculate the total energies of all cells with respect to the CoM by summing up their kinetic, potential, and internal energies. Cells with negative total energies are bound to the initial CoM. In the next iteration, we calculate the new values of $\mathbf{r}_{\mathrm{CoM}}$ and $\mathbf{v}_{\mathrm{CoM}}$ of only the bound cells. This process is repeated until the CoM values do not vary.
@@ 3.1 Overview
During a collision event, the two stars go through multiple close passages and ultimately coalesce. During a merger event, two stars in a binary advance closer to each other due to orbit shrinkage, undergo mass transfer and contact, and eventually coalesce into a single object.
@@
As an overview, we show successive moments of the 10 + 5, $b=0.25$ collision and the 10 + 5 merger in Figures 2 and 3, respectively. All figures display mid-plane slices of three quantities: density $\rho$, passive scalar $K$, and magnetic field $B$, at four different times: well before, just before, just after, and well after coalescence. The middle panels indicate that the cores of both the collision and merger products are predominantly made up of material from the less massive stars ($5\,\mathrm{M}_{\odot}$). In other collision cases with different masses, however, collision products are not completely dominated by the initially less massive star. The mixing in the outer layers of the collision and merger products differs, but since the outer layers of both stars are of similar chemical compositions (see Section 3.2) and contribute less mass compared to the cores, this is not expected to affect the stars’ long-term evolution. Finally, both collision and merger products have significantly higher magnetic field strengths than their initial stars, as seen in the bottom panels. However, it is also clear that the magnetic field amplification in the merger product is considerably larger than in the $b=0.25$ collision product (the color scales are the same). This is also true for other $b=0.25$ collision and merger products. However, the collision simulations with $b=0.50$ exhibit magnetic field amplification akin to mergers (see Section 3.5).
@@
Figure 5: Mid-plane radial profiles of mid-plane angular frequency ratios $\omega/\omega_{\mathrm{crit}}$ profiles of collision (left) and merger (right) products for different masses. Merger products rotate faster than collision products and are closer to break-up velocity.
3.2 Chemical mixing
Collision and merger interactions result in mixing in the stars, transporting helium from the cores to the new envelope and hydrogen from the envelopes to the new core, thus rejuvenating the star (see also J. C. Lombardi et al., 1996; F. R. N. Schneider et al., 2016). Figure 4 shows the profiles of $^{1}\mathrm{H}$ (hydrogen) fractions $X$, respectively, for all collision and merger products. Since all stars are in the main-sequence and have $Z_{\star}=0.02$ (see Section 2), the $^{4}\mathrm{He}$ (helium) fractions $Y$ are inverses of $X$ and follow similar trends. The sorting of stellar layers in the resultant products depends on their relative specific entropies or ’buoyancy’ parameters (e.g., J. C. Lombardi et al., 1996; E. Gaburov et al., 2008): less buoyant material sinks to the new stellar core. For massive stars, the buoyancy parameters depend on densities, gas pressures, and radiation pressures (E. Gaburov et al., 2008).
In general, the cores of merger products are dominated by the initially less massive stars, while the cores of collision products contain some material from the initially more massive stars (see also Section 3.1), and this is reflected in the core hydrogen fractions $X_{\mathrm{core}}$. The largest differences in core compositions are seen in the 8 + 7, 10 + 7 and 8 + 5 cases, where $X_{\mathrm{core}}$ of the merger products are higher by $\sim 3\%$, $\sim 6\%$ and $\sim 10\%$, respectively, than the corresponding collision products. The differences in $X_{\mathrm{core}}$ between collision and merger products in the other cases are within $1\%$ of each other.
The stellar envelopes of all collision and merger products have $X\lesssim 0.70$, the hydrogen fraction for isolated MS stellar envelopes. However, the outermost regions, corresponding to the stellar surfaces, do not show significant hydrogen depletion or, alternatively, helium enrichment.
Previous studies have shown that post–main-sequence evolution, in particular whether a star evolves to a blue or red supergiant, is highly sensitive to the hydrogen profile just above the convective core (e.g., A. Schootemeijer et al., 2019; J. Klencki et al., 2022). Subtle differences between collision and merger products, as seen in our collision and merger models, may therefore lead to distinct post–main-sequence outcomes, which should be further explored using 1D stellar evolution models capable of following their long-term evolution.
3.3 Rotation and break-up
Collision and merger products rotate differentially and at notable fractions of their break-up velocities owing to the transfer of orbital angular momenta of the initial systems of stars to the spin angular momenta of the resulting products. Figure 5 shows the ratios of the angular frequencies $\omega$ to the critical (break-up) angular frequencies $\omega_{\mathrm{crit}}=[Gm(r)/r^{3}]^{0.5}$, where $m(r)$ is the enclosed cylindrical mass in radius $r$, in the mid-planes for all collision and merger products at the end of the simulations. The mid-planes are defined as regions around the collision/merger planes within a distance of $0.5\,\mathrm{R}_{\odot}$. $\omega$ values reach maxima of $\sim 0.5$$0.6\,\omega_{\mathrm{crit}}$ within $b=0.25$ collision products, and $\sim 0.8$$0.9\,\omega_{\mathrm{crit}}$ within $b=0.50$ collision products and merger products.
The large spin angular momenta within merger and collision products have significant implications for their long-term evolution. During thermal contraction of the resulting products, large centrifugal forces can potentially result in substantial mass loss unless they efficiently shed their excess spin angular momenta (A. Sills et al., 2001, 2005). A. Sills et al. (2005) suggested that strong magnetic fields generated during coalescence may lead to loss of angular momentum through magnetic braking by winds (E. J. Weber & L. Davis, 1967) or magnetic disk locking (P. Ghosh & F. K. Lamb, 1978). In addition, significant spin-down might occur during thermal contraction due to internal structural readjustments F. R. N. Schneider et al. (2019). Given these various processes that affect stellar spins, the exact long-term spin evolution remains uncertain and requires further study.
Figure 6: Edge-on snapshots of the 8 + 7, $b=0.25$ collision product (top), 8 + 7, $b=0.50$ collision product (middle) and the 8 + 7 merger product (bottom) $\sim 60\,t_{\mathrm{dyn},r_{\mathrm{tot}}}$ after coalescence. All panels illustrate slices, passing through the centers of mass, perpendicular to the plane of collision/merger. The four columns illustrate the slices of densities $\rho$ (leftmost), azimuthal components of magnetic fields $B_{\phi}$ (center-left), plasma beta parameters $\beta_{\mathrm{mag}}$ (center-right), and radial velocities $v_{\mathrm{rad}}$ (rightmost). The color scale of $B_{\phi}$ is logarithmic when $|B_{\phi}|\gtrsim 10^{2}\,\mathrm{G}$ and linear otherwise. Also depicted are the smoothed contours of $\rho$ and $\beta_{\mathrm{mag}}=1$. The effect of rotational flattening is clearly seen in $\rho$, with the contours of the merger product being peanut-shaped. While the $b=0.25$ collision product shows small-scale reversals in $B_{\phi}$, the $b=0.50$ collision product and the merger product have large-scale order. Lastly, magnetized ($\beta_{\mathrm{mag}}\ll 1$), bipolar axial outflows ($v_{\mathrm{rad}}\gtrsim 300\,\mathrm{km}\,\mathrm{s}^{-1}$) are evident in the merger product but absent in the collision products.
3.4 Internal structures
Collision and merger products are not spherically symmetric because of rotation, but are axisymmetric about their rotation axes. Figure 6 shows the final, edge-on slices of density $\rho$, azimuthal component $B_{\phi}$ of the magnetic field (detailed in Section 3.5), plasma beta parameter $\beta_{\mathrm{mag}}$ (detailed in Section 3.6), and radial velocity $v_{\mathrm{rad}}$ (detailed in Section 3.6) for the 8 + 7, $b=0.25$ collision product (top), the 8 + 7, $b=0.50$ collision product (middle), and the 8 + 7 merger (bottom). The features of collision and merger products for the other masses are qualitatively similar to those in the 8 + 7 cases (see also Sections 3.5 and 3.6).
We notice rotational flattening in the edge-on density contours (leftmost column of Figure 6) of both collision and merger products. However, while the density contours (and pressure contours; not shown) of the collision products are ellipsoidal due to rotation, those of the merger product are peanut-shaped, suggesting the presence of low-density material perpendicular to the merger plane (see also Section 3.6). We see that the $b=0.50$ collision product is more flattened (see also T. Ryu et al., 2025) due to a larger initial orbital angular momentum. Finally, the merger product is more extended in the merger plane, seemingly having a disk-like structure.
3.5 Magnetic field strengths
In ideal MHD (see Section 2), magnetic fields are ‘frozen-in’ the stellar plasmas, with no physical magnetic diffusion [^5] There is, however, numerical diffusion at scales comparable to the sizes of the cells.. During collisions and mergers, most of the drastic amplification in the magnetic field arises from small-scale dynamos (e.g., F. Cattaneo, 1999; C. Federrath, 2016; A. Reiners et al., 2022), where turbulent mixing leads to chaotic stretching, folding, and twisting of magnetic field lines. F. R. N. Schneider et al. (2019) (Extended Data Figure 2) showed that sufficiently resolved simulations of stellar mergers with varying resolutions exhibit similar saturation levels of the total magnetic energy $E_{\mathrm{B}}$, although their (exponential) growth rates vary, lending credence to small-scale dynamos. Additionally, large-scale material mixing and shearing that occurs owing to differential rotation (see Section 3.3) also leads to some amplification of the magnetic field.
[…]

### v2

@@ II.3 Initial conditions
For cells with densities greater than the threshold given above, $K_{1}+K_{2}\approx 1$. This is not true for very low density cells, since the ‘vacuum’ cells initially have $K_{1}=K_{2}=0$. Thus, $K$ quantifies the mixing through the course of each interaction, with $K>0$ (or $K_{1}>K_{2}$) indicating a higher fraction of the heavier star and $K<0$ (or $K_{1}<K_{2}$) indicating a higher fraction of the lighter star.
@@
Figure 2: Mid-plane snapshots of a stellar collision between a $10\,\mathrm{M}_{\odot}$ star and a $5\,\mathrm{M}_{\odot}$ star when $b=0.25$, $10\,\mathrm{Myr}$ into their MS lifetimes. The three rows illustrate slices, in the collision plane, of densities $\rho$ (top), passive scalars $K$ (middle), and magnetic field magnitudes $B$ (bottom). The four columns illustrate these quantities at the times $t$ before collision (first column), just before collision (second column), just after collision (third column), and after $\sim 60\,t_{\mathrm{dyn},r_{\mathrm{tot}}}$ following collision (fourth column). The core of the collision product is dominated by the initially less massive star, while the outer layers are composed of material from both stars in varying fractions. Furthermore, the amplification of the magnetic field is clearly visible across the panels.
@@
II.3.1 Collisions
We assume the orbits in the collision scenario to be parabolic. This is a reasonable assumption since young stellar clusters are observed to have velocity dispersions $\sigma\sim 10$$20\,\mathrm{km}\,\mathrm{s}^{-1}$ (e.g., Gieles et al., 2010), yielding near-parabolic eccentricities [^3] For a hyperbolic orbit, $e=1+r_{\mathrm{p}}v_{\infty}^{2}/GM$, where $r_{\mathrm{p}}$ is the periapsis distance, $v_{\infty}^{2}\sim\sigma$ is the velocity at infinity, and $M$ is the total mass. of $|e-1|\sim 10^{-4}$ for our typical stellar masses and radii. It should be mentioned that collisions are not very frequent in young stellar clusters; nonetheless, we consider these parameters for comparative purposes. We define the collision impact parameter as $b=r_{\mathrm{p}}/r_{\mathrm{tot}}$, where $r_{\mathrm{p}}$ is the closest approach distance between the CoMs of the two stars. In 6 of the 8 collision simulations, we set $b=0.25$. In the other 2 simulations, we set $b=0.50$ to compare the effect of the impact parameter on the collision product. We initially position the relaxed stars $5\,r_{\mathrm{tot}}$ apart prior to the pericenter passage.
II.3.2 Mergers
Unlike collisions, in reality, mergers occur on much longer timescales. For computational feasibility, we speed up the orbit shrinkage phase by adding artificial deceleration terms (similar to Pakmor et al., 2013, 2021). The equations below, which are the same as Equation 2 in Pakmor et al. (2021) [^4] Note that there is an error in that paper, which is corrected in our equation., quantify the antiparallel, azimuthal acceleration terms that result in an inspiral at a constant rate $v_{\mathrm{insp}}$:
@@ II.3.2 Mergers
$\bm{a}_{\star 1}=-\dfrac{m_{\star 2}^{2}\bm{v}_{\star 1}}{v_{\star 1}^{2}}f(r)$ (4a)
$\bm{a}_{\star 2}=-\dfrac{m_{\star 1}^{2}\bm{v}_{\star 2}}{v_{\star 2}^{2}}f(r)$ (4b)
Here, $r$ is the instantaneous distance between the two stars, $\bm{v}_{\star}$ and $\bm{a}_{\star}$ represent the star’s velocity and acceleration vectors, respectively, and $v_{\mathrm{insp}}=10\,\mathrm{km}\,\mathrm{s}^{-1}$ is the inspiral velocity we choose. In comparison, typical Keplerian circular velocities for the masses and distances considered are $\sim 500$$600\,\mathrm{km}\,\mathrm{s}^{-1}$. This deviation from the circular velocity results in a slightly eccentric inspiral with $e\approx 0.1$. For a constant value of $v_{\mathrm{insp}}$, stars with a more skewed mass ratio merge more quickly (see Section III and Figure 8).
We position the relaxed (nonrotating) stars in a circular binary orbit with an initial semimajor axis of $1.5\,r_{\mathrm{tot}}$ and run the simulations with the sped-up inspiral. As a comparison, using the Roche lobe radius approximation formula from Eggleton (1983), Roche lobe filling ensues at $\lesssim 1.31\,r_{\mathrm{tot}}$ for our range of mass ratios (see Table 1).
III Results
In this section, we use the shorthand notation $m_{\star 1}$ + $m_{\star 2}$’ to denote an interaction between stars of masses $m_{\star 1}$ and $m_{\star 2}$ (in units of $\mathrm{M}_{\odot}$). We focus on comparing key properties of the collision and merger products, including chemical mixing, rotation, magnetic fields, and outflows. While these quantities are intrinsically at least 2D, we discuss them using their mass-weighted averages over equipressure surfaces as representative values [^5] MESA handles stellar rotation in 1D in a similar manner, with pressure being the independent coordinate.. In addition to the averages, we indicate the spread in each calculated quantity within the equipressure surfaces by reporting the 15th and 85th percentiles. In the calculation of the averages and spreads, we only consider cells that are bound to the CoM of the new star, which we calculate iteratively as follows. We first find the initial values of the CoM position $\bm{r}_{\mathrm{CoM}}$ and velocity $\bm{v}_{\mathrm{CoM}}$ of all cells. We then calculate the total energies of all cells with respect to the CoM by summing up their kinetic, potential, and internal energies. Cells with negative total energies are bound to the initial CoM. In the next iteration, we calculate the new values of $\bm{r}_{\mathrm{CoM}}$ and $\bm{v}_{\mathrm{CoM}}$ of only the bound cells. This process is repeated until the CoM values do not vary.
@@ III.1 Overview
During a collision event, the two stars go through multiple close passages and ultimately coalesce. During a merger event, two stars in a binary advance closer to each other owing to orbit shrinkage, undergo mass transfer and contact, and eventually coalesce into a single object.
@@
As an overview, we show successive moments of the 10 + 5, $b=0.25$ collision and the 10 + 5 merger in Figures 2 and 3, respectively. All figures display midplane slices of three quantities: density $\rho$, passive scalar $K$, and magnetic field $B$, at four different times: well before, just before, just after, and well after coalescence. The middle panels indicate that the cores of both the collision and merger products are predominantly made up of material from the less massive stars ($5\,\mathrm{M}_{\odot}$). In other collision cases with different masses, however, collision products are not completely dominated by the initially less massive star. The mixing in the outer layers of the collision and merger products differs, but since the outer layers of both stars are of similar chemical compositions (see Section III.2) and contribute less mass compared to the cores, this is not expected to affect the stars’ long-term evolution. Finally, both collision and merger products have significantly higher magnetic field strengths than their initial stars, as seen in the bottom panels. However, it is also clear that the magnetic field amplification in the merger product is considerably larger than in the $b=0.25$ collision product (the color scales are the same). This is also true for other $b=0.25$ collision and merger products. However, the collision simulations with $b=0.50$ exhibit magnetic field amplification akin to mergers (see Section III.5).
Table 2 shows the bound masses $m_{\mathrm{\star,bound}}$, total energies $E_{\mathrm{\star,tot}}$, and spin angular momenta $L_{\mathrm{\star,ax}}$ about the rotation axes of the collision and merger products after $\sim 60\,t_{\mathrm{dyn},r_{\mathrm{tot}}}$. We estimate the initial orbital energies and angular momenta of the mergers by assuming circular orbits with semimajor axes $r_{\mathrm{tot}}$, where we stop the artificial inspiral (see Equation 3). More than $97\%$ of the total masses remain bound to the resulting products. The total energies of collision products and merger products are within $1\%$$3\%$ and $6\%$$13\%$, respectively, of the summations of the initial orbital energies (relatively negligible for collisions) and the total energies of the initial stars. The spin angular momenta of all resulting products are significant fractions ($62\%$$68\%$) of the orbital angular momenta of the initial systems (see also Section III.3).
Figure 5: Profiles of average angular frequencies in the midplanes (see text for definition) $\omega_{\mathrm{mid}}$ as functions of cylindrical radial distances $s$. The left and right panels represent collision and merger products, respectively, of different masses. Both collision and merger products have sharply declining angular velocity profiles.
Figure 6: Profiles of average angular frequency ratios in the midplanes $\omega_{\mathrm{mid}}/\omega_{\mathrm{crit}}$ as functions of cylindrical radial distances $s$. The left and right panels represent collision and merger products, respectively, of different masses. Merger products rotate faster than collision products and are closer to breakup velocity.
III.2 Chemical mixing
Collision and merger interactions result in mixing in the stars, transporting helium from the cores to the new envelope and hydrogen from the envelopes to the new core, thus rejuvenating the star (see also Lombardi et al., 1996; Schneider et al., 2016). Figure 4 shows the profiles of $^{1}\mathrm{H}$ (hydrogen) fractions $X$ for all collision and merger products. Since all stars are in the MS and have $Z_{\star}=0.02$ (see Section II), the $^{4}\mathrm{He}$ (helium) fractions $Y$ are inverses of $X$ and follow similar trends. The sorting of stellar layers in the resultant products depends on their relative entropies (or ’buoyancy’ parameters; e.g., Lombardi et al., 1996; Gaburov et al., 2008, i.e., less buoyant material sinks to the new stellar core. For massive stars, the buoyancy parameters depend on densities, gas pressures, and radiation pressures (Gaburov et al., 2008). For an updated approach to entropy sorting, see Schneider et al. (2025).
In general, the cores of merger products are dominated by the initially less massive stars, while the cores of collision products contain some material from the initially more massive stars (see also Section III.1), and this is reflected in the core hydrogen fractions $X_{\mathrm{core}}$. The largest differences in core compositions are seen in the 8 + 7, 10 + 7, and 8 + 5 cases, where $X_{\mathrm{core}}$ of the merger products are higher by $\sim 3\%$, $\sim 6\%$, and $\sim 10\%$, respectively, than the corresponding collision products. The differences in $X_{\mathrm{core}}$ between collision and merger products in the other cases are within $1\%$ of each other.
The outermost modeled regions of all collision and merger products have $X\sim 0.70$, which is the hydrogen fraction of isolated MS stellar envelopes. Therefore, we conclude that neither collisions nor mergers will show surface hydrogen depletion or, alternatively, helium enhancement.
Previous studies have shown that post–MS evolution, in particular whether a star evolves to a blue or red supergiant, is highly sensitive to the hydrogen profile just above the convective core (e.g., Schootemeijer et al., 2019; Klencki et al., 2022). Subtle differences between collision and merger products, as seen in our collision and merger models, may therefore lead to distinct post–MS outcomes, which should be further explored using 1D stellar evolution models capable of following their long-term evolution.
III.3 Rotation and breakup
Collision and merger products rotate differentially and at notable fractions of their breakup velocities owing to the transfer of orbital angular momenta of the initial systems of stars to the spin angular momenta of the resulting products. Figure 5 shows the average angular frequencies in the midplanes $\omega_{\mathrm{mid}}$ for all collision and merger products at the end of the simulations. All resulting products have steeply decreasing profiles of $\omega_{\mathrm{mid}}$, resulting in significant differential rotation. The midplanes are defined as regions around the orbital planes within a vertical distance of $0.5\,\mathrm{R}_{\odot}$. Figure 6 shows the ratios of $\omega_{\mathrm{mid}}$ to the critical (breakup) angular frequencies $\omega_{\mathrm{crit}}=[Gm(s)/s^{3}]^{0.5}$, where $m(s)$ is the enclosed cylindrical mass within the cylindrical radius $s$, for all collision and merger products at the end of the simulations. The $\omega_{\mathrm{mid}}$ values reach maxima of $\sim 0.5$$0.6\,\omega_{\mathrm{crit}}$ within $b=0.25$ collision products and $\sim 0.8$$0.9\,\omega_{\mathrm{crit}}$ within $b=0.50$ collision products and merger products.
The large spin angular momenta within merger and collision products have significant implications for their long-term evolution. During thermal contraction of the resulting products, large centrifugal forces can potentially result in substantial mass loss unless they efficiently shed their excess spin angular momenta (Sills et al., 2001, 2005). Sills et al. (2005) suggested that strong magnetic fields generated during coalescence may lead to loss of angular momentum through magnetic braking by winds (Weber and Davis, 1967) or magnetic disk locking (Ghosh and Lamb, 1978). In addition, significant spin-down might occur during thermal contraction owing to internal structural readjustments (Schneider et al., 2019). Given these various processes that affect stellar spins, the exact long-term spin evolution remains uncertain and requires further study.
Figure 7: Edge-on snapshots of the 8 + 7, $b=0.25$ collision product (top); 8 + 7, $b=0.50$ collision product (middle); and the 8 + 7 merger product (bottom) $\sim 60\,t_{\mathrm{dyn},r_{\mathrm{tot}}}$ after coalescence. All panels illustrate slices, passing through the CoMs, perpendicular to the orbital plane. The four columns illustrate the slices of densities $\rho$ (first column), azimuthal components of magnetic fields $B_{\phi}$ (second column), plasma beta parameters $\beta_{\mathrm{mag}}$ (third column), and radial velocities $v_{\mathrm{rad}}$ (fourth column). The color scale of $B_{\phi}$ is logarithmic when $|B_{\phi}|\gtrsim 10^{2}\,\mathrm{G}$ and linear otherwise. Also depicted are the smoothed contours of $\rho$ and $\beta_{\mathrm{mag}}=1$. The effect of rotational flattening is clearly seen in $\rho$, with the contours of the merger product being peanut shaped. While the $b=0.25$ collision product shows small-scale reversals in $B_{\phi}$, the $b=0.50$ collision product and the merger product have large-scale order. Lastly, magnetized ($\beta_{\mathrm{mag}}\ll 1$), bipolar axial outflows ($v_{\mathrm{rad}}\gtrsim 300\,\mathrm{km}\,\mathrm{s}^{-1}$) are evident in the merger product but absent in the collision products.
III.4 Stellar structures
Collision and merger products are not spherically symmetric because of rotation, but they are axisymmetric about their rotation axes. Figure 7 shows the final, edge-on slices of density $\rho$, azimuthal component $B_{\phi}$ of the magnetic field (detailed in Section III.5), plasma beta parameter $\beta_{\mathrm{mag}}$ (detailed in Section III.6), and radial velocity $v_{\mathrm{rad}}$ (detailed in Section III.6) for the 8 + 7, $b=0.25$ collision product (top); the 8 + 7, $b=0.50$ collision product (middle); and the 8 + 7 merger (bottom). The features of collision and merger products for the other masses are qualitatively similar to those in the 8 + 7 cases (see also Sections III.5 and III.6).
We notice rotational flattening in the edge-on density contours (leftmost column of Figure 7) of both collision and merger products. However, while the density contours (and pressure contours; not shown) of the collision products are oblate due to rotation, those of the merger product are peanut shaped, suggesting a dearth of material perpendicular to the merger plane (see also Section III.6). We see that the $b=0.50$ collision product is more flattened (see also Ryu et al., 2025) due to a larger initial orbital angular momentum.
Figure 9: Spacetime (‘butterfly’) diagrams, at cylindrical radii $s\simeq 3\,\mathrm{R}_{\odot}$, of the radial components of magnetic field $\overline{B}_{r}$ (volume averaged). The left, middle, and right panels correspond to the 10 + 5, $b=0.25$ collision product; the 10 + 5, $b=0.50$ collision product; and the 10 + 5 merger product respectively. Note that the colorbar scale in the left panel differs from those in the middle and right panels. In the $b=0.50$ collision product and the merger product, the signs of $\overline{B}_{r}$ show notable outward-moving patterns, characteristic of MRI. This is not the case in the $b=0.25$ collision product.
III.5 Magnetic fields
In ideal MHD (see Section II), magnetic fields are ‘frozen-in’ the stellar plasmas, with no physical magnetic diffusion [^6] There is, however, numerical diffusion at scales comparable to the sizes of the cells.. During collisions and mergers, most of the drastic amplification in the magnetic field arises from small-scale dynamos (e.g., Cattaneo, 1999; Federrath, 2016; Reiners et al., 2022), where turbulent mixing leads to chaotic stretching, folding, and twisting of magnetic field lines. Schneider et al. (2019) (Extended Data Figure 2) showed that sufficiently resolved simulations of stellar mergers with varying resolutions exhibit similar saturation levels of the total magnetic energy $E_{\mathrm{B}}$, although their (exponential) growth rates vary, lending credence to small-scale dynamos. Additionally, large-scale mixing and the magnetorotational instability (MRI; Balbus and Hawley, 1991) arising from differential rotation and shear (see Section III.3) also lead to some amplification of the magnetic field.
@@ III.5 Magnetic fields
To understand the source of this magnetic field amplification, we estimate the turbulent energies, $E_{\mathrm{turb}}$, as the sum of non-azimuthal kinetic energies of the bound stellar material, i.e, $E_{\mathrm{turb}}\simeq\sum_{\mathrm{cells}}0.5\,m_{\mathrm{cell}}\,(v_{\mathrm{s,cell}}^{2}+v_{\mathrm{z,cell}}^{2})$, where $v_{\mathrm{s,cell}}$ and $v_{\mathrm{z,cell}}$ are the cylindrical radial and vertical velocities, respectively, of each cell [^7] A more accurate turbulent energy includes deviation from rotation in the azimuthal direction. However, this term is not included for simplicity.. We find that $E_{\mathrm{B}}$ of the 7 + 5, 8 + 5, 8 + 7, and 10 + 5 collisions with $b=0.25$ is a factor of a few lower than $E_{\mathrm{turb}}$, suggesting that their magnetic field amplification is primarily sourced from small-scale dynamos that arise due to turbulence (see also Ohlmann et al., 2025; Schneider et al., 2025). However, in the other collisions and all mergers, $E_{\mathrm{B}}$ is either similar to or a factor of a few higher than $E_{\mathrm{turb}}$, hinting at other mechanisms, such as MRI, in play.
We test this hypothesis of MRI augmenting magnetic field amplification in Figure 9, which presents the space-time (or ‘butterfly’) diagrams (e.g., Brandenburg et al., 1995; Stone et al., 1996) of the radial components of magnetic field $\overline{B}_{r}$ for the 10 + 5 collision and merger products. Such plots are quite useful for diagnosing the differences in magnetic fields (and other quantities) at different vertical heights $z$ over time $t$. To generate space-time diagrams, we first consider cells in a cylindrical region of radius $s\simeq 3\,\mathrm{R}_{\odot}$ around the axis perpendicular to the orbital plane and passing through the CoM, for each simulation snapshot after coalescence. Next, we bin these selected cells based on their heights relative to the orbital plane. Finally, we volume-average the radial components of magnetic fields in each bin and plot this quantity, $\overline{B}_{r}$, as a function of $z$ and $t$.
Figure 9 shows that $\overline{B}_{r}$ changes sign and rises above the midplane over time in the $b=0.50$ collision product and the merger product but not in the $b=0.25$ collision product. These vertically propagating fields may be the result of (large-scale) dynamo waves (Parker, 1955), similar to the $11\,\mathrm{yr}$ Solar sunspot cycle (e.g., Babcock, 1961), which are driven by differential rotation and MRI. In such dynamos, MRI and turbulence generate poloidal fields ($\alpha$ effect), and differential rotation generates toroidal fields ($\Omega$ effect), both of which are necessary for stable magnetic field configurations (Markey and Tayler, 1973; Tayler, 1973; Braithwaite and Spruit, 2004). Interestingly, unlike other $b=0.25$ collision products, this signature of MRI also develops in the 10 + 8 and 10 + 7 collision products $\sim 100\,\mathrm{hr}$ after collision, supporting the larger (and gradually increasing) $E_{\mathrm{B}}$ in these two cases.